OpenStudy (anonymous):

What is the equation of the circle with center (2, -5) that passes through the point (–2, 10)?

OpenStudy (jdoe0001):

do you know what the "circle formula" is?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

(x - h)^2 + (y - k)^2 =r^2 So, (2 - -2)^2 + (-5 - 10)^2?

OpenStudy (jdoe0001):

ok, well, you're given the (h,k) for the center, which are the "x" and "y" coordinates, which is your 1st point up there now, you just need the radius you have a 2nd point, from the center to a point in the circle, that'd give you the radius since you have 2 points, just use the distance formula to get the distance from THAT point to THE center point, and that'd give you the radius http://0.tqn.com/d/math/1/0/A/1/distance.gif

OpenStudy (jdoe0001):

well, yeah, your way also works :)

OpenStudy (anonymous):

(x – (–2))^2 + (y – 10)^2 = 241